# Crossing the Threshold: Idiomatic Machine Translation through Retrieval Augmentation and Loss Weighting

### v1

@@ 3 Idioms and Data Collection
Finally, we collect a random test set from an alternate source, the Ted Talks corpus Reimers and Gurevych (2020). This is to ensure that translation quality of other, unrelated sentences is not impacted by any modifications meant to improve translation of idioms. We collect sentences from Ted Talks, rather than OpenSubtitles, because it allows us to also examine a domain shift. This is because topics discussed and vocabulary used in Ted Talks may be slightly different from what is discussed in movies or TV shows. To control for translation length as a source of difficulty, sentences were length-matched on the target side with corresponding sentences in the idiomatic set. This created the random set, which is the same size as the idiomatic test set. All three test sets are summarized in Table 2.
| Language | Idiom matches | Idiomatic | Literal | Random | Total |
| fr | 85632 | 777 | 89 | 777 | 1643 |
| fi | 51811 | 449 | 84 | 449 | 982 |
| ja | 23018 | 3253 | 389 | 3253 | 6895 |
@@ 4.2 Evaluation of Commercial Systems
Although it’s impossible for us to determine what data these commercial systems were trained on, we examine the frequency of each idiom within OpenSubtitles as a proxy for its overall frequency in the training data, and bucket idioms into quintiles based on their occurrence frequency in source text. As idioms become more frequent, the quality of translations increases. An example of DeepL on the French idiom set is shown in Figure 2. Trends for other languages and systems are in Appendix G. This indicates that like in the synthetic experiments, there may be strong frequency effects on translation quality of idioms.
@@ 6.3 Evaluation
We use multiple automatic metrics to evaluate translation quality. However, due to the importance of accurate semantic evaluation, the authors (native English speakers and fluent in French and Japanese) conduct a human evaluation inspired by MQM Lommel et al. (2014). Only errors that would fall under the“terminology” and “accuracy” error types are considered, as we are focused on severe semantic errors. We give a score of 0 for severe errors and a score of 0.5 for major errors. A score of 1 is given otherwise. Exact evaluation standards are in Appendix D.
@@ 7.1 Automatic and Human Evaluation
In most cases, as reported in Figure 3, using a combination of sentence upweighting and kNN-MT led to the greatest increase in automatic metrics on all three test sets, of up to 3.08 BLEU points on the idiomatic test set (fr), 2.69 BLEU points on the literal test set (fi), and 5.75 points on the random test set (fr). In all cases except ja-rand, using one or more of these methods improved over the baseline. Exact numerical results are in Appendix I.
We evaluate the statistical significance of the results through a one-tailed permutation test Graham et al. (2014). Further details are in Appendix E. Exact results are in Appendix F. For Finnish, significance is achieved for all three test sets, and for French, significance is achieved for the idiomatic and random test sets. For Japanese, values achieved are not significant, but are borderline.
@@ 7.2 Error analysis
We repeat the frequency analysis performed on commercial systems () for LM, and find that adding upweighting and kNN-MT generally improves translations at all frequency levels. These increases are not concentrated in low-frequency idioms, so more common idioms continue to be translated better.888This trend is different in retrieval of long-tail facts in question answering, in which retrieval flattens out the difference between rare and common facts Kandpal et al. (2022). A representative example (for French) is in Figure 4. A complete set of plots are in Appendix H.
@@ Acknowledgements
Thank you to Perez Ogayo for helping with DeltaLM setup as well as to NeuLab members for providing feedback on parts of the draft. This project was funded by the P2020 program MAIA (LISBOA-01-0247-FEDER-045909).
@@ Appendix C OpenSubtitles Training Details
Appendix D Standards for Human Evaluation
@@
Appendix E Statistical Significance Testing
@@
Appendix F Statistical Significance Results
@@
Appendix G Effect of Frequency on Commercial Translations
Appendix H Effect of Frequency on LM
Appendix I Automatic Metrics in Detail
@@
| Random | Normal | 0.2365 | 0.535 | 0.9145 | 0.4971 | |
@@
| Random | Normal | 0.2778 | 0.528 | 0.9377 | 0.5504 | |
@@
| Random | Normal | 0.0948 | 0.3539 | 0.8946 | 0.3436 | |

### v2

@@ 3 Idioms and Data Collection
To validate these judgments, we hired native annotators in French and Finnish. They were presented with examples from the final literal and idiomatic test sets in a shuffled order, and asked to label them with idiomatic, literal, or N/A labels if they didn’t think it was an instance of either. Agreement (Krippendorff’s Krippendorff (1970); Castro (2017)) in both cases was moderately high (French , Finnish ). Details can be found in Appendix D.
Finally, we collect two random test sets, one which is in-domain and another which is out-of-domain. For the in-domain test set, we simply select sentences from the development set of OpenSubtitles (see subsection 6.2 for details on our split of OpenSubtitles). For the out-of-domain test set, we use the Ted Talks corpus Reimers and Gurevych (2020). This is to ensure that translation quality of other, unrelated sentences is not impacted by any modifications meant to improve translation of idioms. Topics discussed and vocabulary used in Ted Talks may be slightly different from what is discussed in movies or TV shows, so training the model on OpenSubtitles and testing on Ted Talks allows us to evaluate model generalization. For both test sets, to control for translation length as a source of difficulty, sentences were length-matched on the target side with corresponding sentences in the idiomatic set. This created the random set, which is the same size as the idiomatic test set. All three test sets are summarized in Table 2.
| Language | Idiom matches | Idiomatic | Literal | Random (in) | Random (out) | Total |
| fr | 85632 | 777 | 79 | 777 | 777 | 2410 |
| fi | 51811 | 449 | 81 | 449 | 449 | 1428 |
| ja | 23018 | 3253 | 389 | 3253 | 3253 | 10148 |
@@ 4.2 Evaluation of Commercial Systems
Although it’s impossible for us to determine what data these commercial systems were trained on, we examine the frequency of each idiom within OpenSubtitles as a proxy for its overall frequency in the training data, and bucket idioms into quintiles based on their occurrence frequency in source text. As idioms become more frequent, the quality of translations increases. An example of DeepL on the French idiom set is shown in Figure 2. Trends for other languages and systems are in Appendix H. This indicates that like in the synthetic experiments, there may be strong frequency effects on translation quality of idioms.
@@ 6.3 Evaluation
We use multiple automatic metrics to evaluate translation quality. However, due to the importance of accurate semantic evaluation, the authors (native English speakers and fluent in French and Japanese) conduct a human evaluation inspired by MQM Lommel et al. (2014). Only errors that would fall under the“terminology” and “accuracy” error types are considered, as we are focused on severe semantic errors. We give a score of 0 for severe errors and a score of 0.5 for major errors. A score of 1 is given otherwise. Exact evaluation standards are in Appendix E.
@@ 7.1 Automatic and Human Evaluation
In most cases, as reported in Figure 3, using a combination of sentence upweighting and kNN-MT led to the greatest increase in automatic metrics on all three test sets, of up to 3.08 BLEU points on the idiomatic test set (fr), 2.69 BLEU points on the literal test set (fi), and 5.75 points on the random test set (fr). In all cases except ja-rand, using one or more of these methods improved over the baseline. Exact numerical results are in Appendix J.
We evaluate the statistical significance of the results through a one-tailed permutation test Graham et al. (2014). Further details are in Appendix F. Exact results are in Appendix G. For Finnish, significance is achieved for all three test sets, and for French, significance is achieved for the idiomatic and random test sets. For Japanese, values achieved are not significant, but are borderline.
@@ 7.2 Error analysis
We repeat the frequency analysis performed on commercial systems () for LM, and find that adding upweighting and kNN-MT generally improves translations at all frequency levels. These increases are not concentrated in low-frequency idioms, so more common idioms continue to be translated better.888This trend is different in retrieval of long-tail facts in question answering, in which retrieval flattens out the difference between rare and common facts Kandpal et al. (2022). A representative example (for French) is in Figure 4. A complete set of plots are in Appendix I.
@@ Acknowledgements
Thank you to Perez Ogayo for helping with DeltaLM setup, all annotators who validated idiomatic/literal judgments, as well as to NeuLab members for providing feedback on parts of the draft. This project was funded by the P2020 program MAIA (LISBOA-01-0247-FEDER-045909).
@@ Appendix C OpenSubtitles Training Details
Appendix D Native Annotator Recruiting
Annotators were hired through Upwork, and consisted of professional translators in French and Finnish who were also fluent in English. Annotators were paid between USD an hour, depending on their individual hourly rate. No personally identifying information was collected.
Annotators were shown the task description below, and also had access to a compiled list of literal and figurative translations for each idiom.
Description:
This job is to create a dataset that will allow us (researchers at Carnegie Mellon University, PI Graham Neubig) to study the ability of machine translation systems to translate idioms. We would like you to look at translations and decide whether they are literal or idiomatic translations. These will be English sentences containing the translation of a sentence in either French, Finnish, or Japanese. All the sentences contain a phrase that is an idiom in the foreign language, but sometimes the translations may be meant literally. As an example, in English “kick the bucket” can mean to die when used in an idiomatic way, or to literally kick a bucket when used in a literal way. You should mark the example with “idiomatic” if you think it’s an idiomatic example, or “literal” if you think it’s a literal example. If you think the translation is not a good translation, we will ask you to mark this entry with the word “None”, but this should occur in relatively few cases. We estimate that the job will take 2-3 hours overall. Examples: Original sentence: Jalkaväkeä oli pilvin pimein. English translation (reference): Arty and infantry just kept on coming. Contains idiom: pilvin pimein Idiom literal meaning: In dark clouds Idiom figurative meaning: A huge (often excessive) amount of something. Label: idiomatic Original sentence: Harmi, ettei lehdistötilaisuudessasi ollut lammasta. English translation (reference): Too bad they didn’t have lamb at your press conference, Francis. Contains idiom: olla lammas Idiom literal meaning: To be a lamb Idiom figurative meaning: a person who is like a lamb does nothing alone. the person does Label: literal
Appendix E Standards for Human Evaluation
@@
Appendix F Statistical Significance Testing
@@
Appendix G Statistical Significance Results
@@
Appendix H Effect of Frequency on Commercial Translations
Appendix I Effect of Frequency on LM
Appendix J Automatic Metrics in Detail
@@ Appendix I Effect of Frequency on LM
| Random-in | Normal | 0.2056 | 0.5044 | 0.9321 | 0.4739 | |
| Knn | 0.1991 | 0.5067 | 0.9316 | 0.4732 | ||
| Upweight | 0.2044 | 0.5048 | 0.9324 | 0.4728 | ||
| Knn + upweight | 0.2120 | 0.5224 | 0.9351 | 0.488 | ||
| Random-out | Normal | 0.2365 | 0.535 | 0.9145 | 0.4971 | |
@@
| Random-in | Normal | 0.3197 | 0.5942 | 0.9443 | 0.5803 | |
| Knn | 0.3142 | 0.5949 | 0.9438 | 0.5774 | ||
| Upweight | 0.3079 | 0.5884 | 0.9427 | 0.5715 | ||
| Knn + upweight | 0.3258 | 0.6036 | 0.9461 | 0.5865 | ||
| Random-out | Normal | 0.2778 | 0.528 | 0.9377 | 0.5504 | |
@@
| Random-in | Normal | 0.1443 | 0.4008 | 0.9212 | 0.3927 | |
| Knn | 0.1535 | 0.4036 | 0.9218 | 0.3960 | ||
| Upweight | 0.1430 | 0.3981 | 0.9212 | 0.3925 | ||
| Knn + upweight | 0.1543 | 0.407 | 0.9224 | 0.4009 | ||
| Random-out | Normal | 0.0948 | 0.3539 | 0.8946 | 0.3436 | |